OpenStudy (anonymous):

place the answer in scentific notation, when appropriate and simplify the units.

OpenStudy (anonymous):

\[T _{p}=2\pi \sqrt{\frac{ 125.4 cm }{ 9.81 m/s ^{2}}}\]

OpenStudy (fifciol):

convert 125.4 cm to meters

OpenStudy (anonymous):

1.254 m

OpenStudy (fifciol):

yes, now just calculate

OpenStudy (anonymous):

what about the m/s^2?

OpenStudy (fifciol):

this is in SI units you don't have to convert it

OpenStudy (anonymous):

oh ok. thanks :)

OpenStudy (fifciol):

yw:)

OpenStudy (anonymous):

@Fifciol will the answer be 2.24 m?

OpenStudy (fifciol):

Period is in seconds so it should be: T=2,246 s

OpenStudy (anonymous):

oh ok :) thanks